OpenStudy (anonymous):

Find the derivative using rules. (x^2+3)/(x-1)^3+(x+1)^3

OpenStudy (callisto):

Is the question (i) \(\frac{(x^2+3)}{(x-1)^3+(x+1)^3}\) or (ii) \(\frac{(x^2+3)}{(x-1)^3}+(x+1)^3\)?

OpenStudy (anonymous):

First one

OpenStudy (callisto):

Okay. Do you know quotient rule and chain rule?

OpenStudy (anonymous):

quotient rule yes, we have not learned the chain rule yet...

OpenStudy (callisto):

Do you know how to differentiate \((x-1)^3\)?

OpenStudy (anonymous):

just use power rule lol

OpenStudy (anonymous):

\[3(x-1)^{2}\]?

OpenStudy (anonymous):

^

OpenStudy (callisto):

Yup. What's (x-1)^3 + (x+1)^3?

OpenStudy (anonymous):

3(x−1)^2 + 3(x+1)^2

OpenStudy (callisto):

And (x^2+3) ?

OpenStudy (anonymous):

2x

OpenStudy (callisto):

Ok. Then you're ready to apply quotient rule! \[\frac{d}{dx}(\frac{u}{v}) = \frac{vu'-uv'}{v^2}\] Let say, in your question, u = x^2+3, what is v?

OpenStudy (anonymous):

.3(x−1)^2 + 3(x+1)^2? o_o

OpenStudy (anonymous):

lol

OpenStudy (callisto):

No o_o

OpenStudy (callisto):

Compare \(\huge\frac{u}{v}\) and \(\huge \frac{(x^2+3)}{(x-1)^3+(x+1)^3}\) If u = x^2+3 what is v?

OpenStudy (anonymous):

The denominator...(x-1)^3 + (x+1)^3

OpenStudy (callisto):

Yes!!!!! Now, u' is the derivative of u(with respect to x); v' is the derivative of v (with respect to x). u =x^2+3 v = (x-1)^3 + (x+1)^3 u' = ??? v' = ???

OpenStudy (anonymous):

u = 2x v = 6(x^2+1)?

OpenStudy (anonymous):

u' v'. Correction :)

OpenStudy (callisto):

Noooooo. Check v' again!!

OpenStudy (callisto):

Finding v' is the same as finding the derivative of (x-1)^3 + (x+1)^3

OpenStudy (anonymous):

I'm lost then. Why isn't it 3(x−1)^2 + 3(x+1)^2

OpenStudy (callisto):

It is. u =x^2+3 u' = 2x v = (x-1)^3 + (x+1)^3 v' = 3(x−1)^2 + 3(x+1)^2 Now, plug everything into \[\frac{d}{dx}(\frac{u}{v}) = \frac{vu'-uv'}{v^2}\] Then you can get your answer.

OpenStudy (anonymous):

Great, thanks!

OpenStudy (callisto):

Welcome~